# MQFL-FHE: Multimodal Quantum Federated Learning Framework with Fully Homomorphic Encryption

### v1

@@ II Related Works
Privacy preservation in FL: (Rahulamathavan et al., 2023) proposed FheFL, which incorporates a multi-key additive homomorphic encryption scheme to mitigate privacy risks associated with gradient sharing in FL. The experimental results confirmed that FheFL maintains model accuracy while providing robust security against potential data breaches, emphasizing the necessity of integrating FHE techniques into federated learning frameworks to enhance data privacy. FedSHE was created by (Pan et al., 2024). It is a useful FL scheme that uses adaptive segmented CKKS homomorphic encryption to keep gradients safe from leakage attacks. Their analysis of security parameters underscores the trade-offs between correctness, efficiency, and security in FL environments.
@@ IV-A QFL-FHE: Integrating Fully Homomorphic Encryption with Quantum Federated Learning
The CKKS (Cheon-Kim-Kim-Song) encryption scheme is particularly well-suited for federated learning because it allows approximate arithmetic on encrypted data, enabling repeated additions and multiplications while maintaining fidelity. Each client $i$ holds local data $x_{i}$ and computes model updates through a local update function $g_{i}$. The encrypted model updates can be expressed as:
$Enc_{\text{CKKS}}(g_{i}(\theta,x_{i}))=\sum_{j=0}^{N}c_{j}\cdot X^{j}+E,$
@@
$Enc_{\text{CKKS}}\left(\sum_{i=1}^{M}g_{i}(\theta,x_{i})\right)=\sum_{j=0}^{N}c_{j}\cdot\left(\sum_{i=1}^{M}g_{i}(\theta,x_{i})\right)^{j}+E^{\prime},$
@@
$|x_{i}\rangle\equiv U^{\dagger}(x_{i})|0\rangle=\bigotimes_{j=1}^{dh}R_{x}^{\dagger}(x_{ij})|0\rangle,$
@@
$C_{\text{global}}=\text{Enc}_{\text{CKKS}}\left(\frac{1}{N}\sum_{i=1}^{N}M(\boldsymbol{\theta})\right)=\frac{1}{N}\sum_{i=1}^{N}C_{i}.$
@@
$|y\rangle=M(\boldsymbol{\theta})|x\rangle=U_{k}(\theta_{k})\cdots U_{1}(\theta_{1})|x\rangle,$
@@
$M(\theta)=\prod_{l=1}^{L}\left(\prod_{i=1}^{n}R_{X}(\theta_{l,i})\right)\left(\prod_{i=1}^{n-1}\text{CNOT}(i,i+1)\right),$
@@ IV-B MQFL-FHE: Multimodal Quantum Federated Learning with Fully Homomorphic Encryption
$\mathbf{w}_{k}=M_{k}(\mathbf{w}_{g},\mathcal{D}_{k})=M_{k}\left(\mathbf{w}_{g},\bigcup_{m=1}^{M}\mathcal{D}_{k,m}\right)$
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$
@@
$S=\sum_{k=1}^{N}\left(\mathbf{w}_{k}^{enc}\cdot\frac{n_{k}}{n_{\text{total}}}\right)=\sum_{k=1}^{N}\left(\text{Enc}_{\text{CKKS}}\left(M_{k}(\boldsymbol{\theta},\mathcal{D}_{k})\right)\cdot\frac{n_{k}}{n_{\text{total}}}\right),$
@@
$g_{k}(x)=\text{softmax}(W_{g}\cdot\phi(x))\quad\text{and}\quad\mathbf{y}_{k}=\sum_{j=1}^{m}g_{k,j}\cdot e_{k,j}(x),$
@@ V Discussion
Figure 4: ROC curves for DNA and MRI datasets under FL with and without quantum enhancements. Panels (a) and (b) show the ROC curves for the DNA dataset under FL-FHE and QFL-FHE setups, respectively, highlighting the performance across various classes. Panels (c) and (d) show the ROC curves for the MRI dataset under FL-FHE and QFL-FHE setups, respectively. Each panel demonstrates the true positive rate against the false positive rate for each class, including micro-average and macro-average ROC curves, illustrating the models’ discriminative ability in a privacy-preserving federated learning context.
In our study, we extensively analyze the performance of federated learning models applied to DNA and MRI datasets, with and without the integration of quantum computing enhancements and FHE. Fig.4 presents these models’ receiver operating characteristic (ROC) curves, offering a detailed view of their discriminative abilities under different configurations. For the DNA dataset, the ROC analysis reveals that the classical FL-FHE model achieves a micro-average and macro-average area under the curve (AUC) of 0.92, indicating robust predictive performance across multiple classes. However, when quantum enhancements are applied through the QFL-FHE setup, there is a noticeable improvement, with the micro-average AUC increasing to 0.94 and the macro-average rising to 0.95. This enhancement validates that integrating QC can refine the model’s sensitivity and specificity, particularly in managing complex patterns within the DNA sequences. In contrast, the MRI dataset exhibits different dynamics. Under the FL-FHE setup, the micro-average and macro-average ROC AUCs are reported at 0.93, demonstrating high efficacy in class discrimination. In the QFL-FHE setup, these metrics improve to 0.97 for both micro-average and macro-average AUCs. This significant increase underscores the potential of quantum enhancements in effectively handling the intricacies of MRI data, potentially due to the quantum model’s capability to process high-dimensional data more efficiently.
@@ V-A Ablation Study
  • Centralized systems outperform federated ones, but the trend varies with QC: Notably, analysis reveals that centralized systems generally outperform federated ones in terms of computational efficiency and ease of implementing privacy-preserving techniques. However, federated systems using QML can surpass centralized models under certain conditions. Specifically, with smaller datasets, QML in federated learning performs better due to its ability to generate richer representations, which is particularly effective for smaller problems. In federated settings, where data is split across clients, QML’s strengths in handling sparse and fragmented data lead to improved performance for decentralized problems, making it more effective than centralized models for such datasets.
@@ VI Conclusion
Our study has shown that FL, despite its advantages for collaborative model training across decentralized data sources, faces significant challenges in data privacy and computational efficiency. These challenges are particularly critical when high accuracy and quick processing are required, such as in real-time applications. The introduction of FHE has proven effective in enhancing data privacy by allowing computations on encrypted data. However, FHE increases the computational load significantly, which can reduce the performance and speed of model training. To tackle these issues, our research introduces an integration of QC with FL and FHE. This new approach has shown potential in reducing FHE’s performance drawbacks and efficiently managing complex, multimodal datasets. Our MQMoE framework demonstrates how QC can improve the learning performance and privacy of federated learning systems. The results from various experimental setups, including classical and quantum models, indicate that while quantum solutions add complexity, they offer potential computational advantages. These findings suggest that quantum-enhanced federated learning could be a viable solution for handling large-scale, privacy-sensitive applications. In summary, this research points out the limitations of current FL systems and proposes a novel solution that uses QC to enhance security and efficiency. While QC is not without its challenges, it represents a promising start in the search for more effective FL technologies. This could be particularly valuable in fields where privacy and quick data processing are essential, marking a significant step forward in developing advanced FL technologies.
@@ VII-A Mathematical Analysis of Error Propagation and Noise Stabilization in FHE-Quantum Hybrid Models
$\Delta_{az}(t)=\Delta_{az}(\vec{v}\cdot SP(t),\vec{v}_{err}\cdot SP(t)),$
$\Delta_{el}(t)=\Delta_{el}(\vec{v}\cdot SP(t),\vec{v}_{err}\cdot SP(t)),$
@@
$J=\begin{bmatrix}0&-\phi-\psi&\theta\\ \phi+\psi&0&0\\ -\theta&0&0\end{bmatrix}.$
@@
$\omega=\frac{2\pi}{\sqrt{\theta^{2}+(\phi+\psi)^{2}}}.$
@@
  • Quantum Noise Confinement: Quantum noise, unlike classical noise, evolves under the unitary constraints of $SU(2)$. The preservation of the norm ensures that errors remain bounded, preventing indefinite accumulation. The bounded evolution introduces oscillatory noise patterns, where errors are periodically ”reset” to smaller magnitudes.
  • Error Correction Through Hybridization: The output of the quantum layer, with its stabilized noise, is passed to the classical neural network. The classical network processes the signal while suppressing residual quantum noise, leveraging the strengths of both paradigms to enhance overall error tolerance(Kundu and Ghosh [2024])
@@
$SP(t)=\exp(tJ),\quad\text{where }J\text{ is traceless and governs periodicity}.$
@@ VII-B Effects of FHE Modifications
Bit Scale Polymodulus Degree Extrema Count Encrypted Parameter Count Time (s) 40 32768 8 15000 290 30 16384 8 15000 17 20 8192 8 20000 319 20 8192 1 50000 209 30 8192 1 85000 338 40 4096 1 180000 440 40 8192 1 140000 326 30 8192 1 350000 391 20 4096 1 410000 339 20 32768 1 690000 432
The main insight from the above table is that reducing parameters like the bit scale, polynomial modulus degree, or extrema count enables encrypting more values but often comes at the expense of security. For example, decreasing the bit scale from $40$ to $20$ increases the number of encryptable values to $690,000$, but it also lowers the security level. Similarly, a smaller polynomial modulus degree allows for encrypting more values, but it reduces both the security level and the depth of operations. In FHE systems, communication costs are as critical as computational efficiency, especially when encrypted data is transmitted, such as in cloud-based machine learning scenarios.
Larger ciphertexts, which result from higher encryption parameters or deeper computational models, require greater bandwidth, increasing communication overhead. Thus, optimizing encryption parameters like bit scale or polynomial degree is crucial for balancing security, computational efficiency, and communication costs.
@@ VII-C Efficacy of MQMoE Approach
The dataset by Livingstone and Russo [2018] contains audio-visual recordings of 24 professional actors (12 male and 12 female), each expressing eight emotions: calm, happy, sad, angry, fearful, surprised, disgust, and neutral. The data is synchronized with both speech and song modalities, making it ideal for multimodal emotion recognition tasks that require the joint analysis of audio and visual cues. This richness in data supports the evaluation of the MQMoE approach in handling real-world, emotionally nuanced inputs across different modalities. The findings from the RAVDESS dataset indicate that the MQMoE framework replicates trends observed in prior experiments. Specifically, the application of quantum federated learning and homomorphic encryption (QFL+FHE) combined yields enhanced performance metrics. The incorporation of audio-visual modalities improves recognition accuracy and proves the modularity and adaptability inherent in the MQMoE architecture. This demonstrates its effectiveness in multimodal emotion recognition, presenting a solution that can adapt to multiple modalities.

### v2

@@ II Related Works
Privacy preservation in FL: (Rahulamathavan et al., 2023) proposed FheFL, which incorporates a multi-key additive homomorphic encryption scheme to mitigate privacy risks associated with gradient sharing in FL. The experimental results confirmed that FheFL maintains model accuracy while providing robust security against potential data breaches, emphasizing the necessity of integrating FHE techniques into FL frameworks to enhance data privacy. FedSHE was created by (Pan et al., 2024). It is a useful FL scheme that uses adaptive segmented CKKS homomorphic encryption to keep gradients safe from leakage attacks. Their analysis of security parameters underscores the trade-offs between correctness, efficiency, and security in FL environments.
@@ IV-A QFL-FHE: Integrating Fully Homomorphic Encryption with Quantum Federated Learning
The Cheon-Kim-Kim-Song (CKKS) encryption scheme is particularly well-suited for FL because it allows approximate arithmetic on encrypted data, enabling repeated additions and multiplications while maintaining fidelity. Each client $i$ holds local data $x_{i}$ and computes model updates through a local update function $g_{i}$. The encrypted model updates can be expressed as:
$Enc_{\text{CKKS}}(g_{i}(\theta,x_{i}))=\sum_{j=0}^{N}c_{j}\cdot X^{j}+E,$ (1)
@@
$Enc_{\text{CKKS}}\left(\sum_{i=1}^{M}g_{i}(\theta,x_{i})\right)=\sum_{j=0}^{N}c_{j}\cdot\left(\sum_{i=1}^{M}g_{i}(\theta,x_{i})\right)^{j}+E^{\prime},$ (2)
@@
$|x_{i}\rangle\equiv U^{\dagger}(x_{i})|0\rangle=\bigotimes_{j=1}^{dh}R_{x}^{\dagger}(x_{ij})|0\rangle,$ (3)
@@
$C_{\text{global}}=\text{Enc}_{\text{CKKS}}\left(\frac{1}{N}\sum_{i=1}^{N}M(\boldsymbol{\theta})\right)=\frac{1}{N}\sum_{i=1}^{N}C_{i}.$ (4)
@@
$|y\rangle=M(\boldsymbol{\theta})|x\rangle=U_{k}(\theta_{k})\cdots U_{1}(\theta_{1})|x\rangle,$ (5)
@@
$M(\theta)=\prod_{l=1}^{L}\left(\prod_{i=1}^{n}R_{X}(\theta_{l,i})\right)\left(\prod_{i=1}^{n-1}\text{CNOT}(i,i+1)\right),$ (6)
@@ IV-B MQFL-FHE: Multimodal Quantum Federated Learning with Fully Homomorphic Encryption
$\mathbf{w}_{k}=M_{k}(\mathbf{w}_{g},\mathcal{D}_{k})=M_{k}\left(\mathbf{w}_{g},\bigcup_{m=1}^{M}\mathcal{D}_{k,m}\right),$ (7)
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (8)
@@
$S$ $=\sum_{k=1}^{N}\left(\mathbf{w}_{k}^{enc}\cdot\frac{n_{k}}{n_{\text{total}}}\right)$
$=\sum_{k=1}^{N}\left(\text{Enc}_{\text{CKKS}}\left(M_{k}(\boldsymbol{\theta},\mathcal{D}_{k})\right)\cdot\frac{n_{k}}{n_{\text{total}}}\right),$ (9)
@@
$g_{k}(x)$ $=\text{softmax}(W_{g}\cdot\phi(x))\quad\text{and}$
$\mathbf{y}_{k}$ $=\sum_{j=1}^{m}g_{k,j}\cdot e_{k,j}(x),$ (10)
@@ V Discussion
Figure 4: ROC curves for DNA and MRI datasets under FL with and without quantum enhancements. Panels (a) and (b) show the ROC curves for the DNA dataset under FL-FHE and QFL-FHE setups, respectively, highlighting the performance across various classes. Panels (c) and (d) show the ROC curves for the MRI dataset under FL-FHE and QFL-FHE setups, respectively. Each panel demonstrates the true positive rate against the false positive rate for each class, including micro-average and macro-average ROC curves, illustrating the models’ discriminative ability in a privacy-preserving FL context.
In our study, we extensively analyze the performance of FL models applied to DNA and MRI datasets, with and without the integration of quantum computing enhancements and FHE. Fig. 4 presents these models’ receiver operating characteristic (ROC) curves, offering a detailed view of their discriminative abilities under different configurations. For the DNA dataset, the ROC analysis reveals that the classical FL-FHE model achieves a micro-average and macro-average area under the curve (AUC) of 0.92, indicating robust predictive performance across multiple classes. However, when quantum enhancements are applied through the QFL-FHE setup, there is a noticeable improvement, with the micro-average AUC increasing to 0.94 and the macro-average rising to 0.95. This enhancement validates that integrating QC can refine the model’s sensitivity and specificity, particularly in managing complex patterns within the DNA sequences. In contrast, the MRI dataset exhibits different dynamics. Under the FL-FHE setup, the micro-average and macro-average ROC AUCs are reported at 0.93, demonstrating high efficacy in class discrimination. In the QFL-FHE setup, these metrics improve to 0.97 for both micro-average and macro-average AUCs. This significant increase underscores the potential of quantum enhancements in effectively handling the intricacies of MRI data, potentially due to the quantum model’s capability to process high-dimensional data more efficiently.
@@ V-A Ablation Study
  • Centralized systems outperform federated ones, but the trend varies with QC: Notably, analysis reveals that centralized systems generally outperform federated ones in terms of computational efficiency and ease of implementing privacy-preserving techniques. However, federated systems using QML can surpass centralized models under certain conditions. Specifically, with smaller datasets, QML in FL performs better due to its ability to generate richer representations, which is particularly effective for smaller problems. In federated settings, where data is split across clients, QML’s strengths in handling sparse and fragmented data lead to improved performance for decentralized problems, making it more effective than centralized models for such datasets.
@@ VI Conclusion
Our study has shown that FL, despite its advantages for collaborative model training across decentralized data sources, faces significant challenges in data privacy and computational efficiency. These challenges are particularly critical when high accuracy and quick processing are required, such as in real-time applications. The introduction of FHE has proven effective in enhancing data privacy by allowing computations on encrypted data. However, FHE increases the computational load significantly, which can reduce the performance and speed of model training. To tackle these issues, our research introduces an integration of QC with FL and FHE. This new approach has shown potential in reducing FHE’s performance drawbacks and efficiently managing complex, multimodal datasets. Our MQMoE framework demonstrates how QC can improve the learning performance and privacy of FL systems. The results from various experimental setups, including classical and quantum models, indicate that while quantum solutions add complexity, they offer potential computational advantages. These findings suggest that quantum-enhanced FL could be a viable solution for handling large-scale, privacy-sensitive applications. In summary, this research points out the limitations of current FL systems and proposes a novel solution that uses QC to enhance security and efficiency. While QC is not without its challenges, it represents a promising start in the search for more effective FL technologies. This could be particularly valuable in fields where privacy and quick data processing are essential, marking a significant step forward in developing advanced FL technologies.
@@ VII-A Mathematical Analysis of Error Propagation and Noise Stabilization in FHE-Quantum Hybrid Models
$\Delta_{az}(t)=\Delta_{az}(\vec{v}\cdot SP(t),\vec{v}_{err}\cdot SP(t)),$ (13)
$\Delta_{el}(t)=\Delta_{el}(\vec{v}\cdot SP(t),\vec{v}_{err}\cdot SP(t)),$ (14)
@@
$J=\begin{bmatrix}0&-\phi-\psi&\theta\\ \phi+\psi&0&0\\ -\theta&0&0\end{bmatrix}.$ (15)
@@
$\omega=\frac{2\pi}{\sqrt{\theta^{2}+(\phi+\psi)^{2}}}.$ (16)
@@
  • Quantum Noise Confinement: Quantum noise, unlike classical noise, evolves under the unitary constraints of $SU(2)$. The preservation of the norm ensures that errors remain bounded, preventing indefinite accumulation. The bounded evolution introduces oscillatory noise patterns, where errors are periodically “reset” to smaller magnitudes.
  • Error Correction Through Hybridization: The output of the quantum layer, with its stabilized noise, is passed to the classical neural network. The classical network processes the signal while suppressing residual quantum noise, leveraging the strengths of both paradigms to enhance overall error tolerance (Kundu and Ghosh [2024]).
@@
$SP(t)=\exp(tJ),\quad\text{where }J\text{ is traceless and governs periodicity}.$ (17)
@@ VII-B Effects of FHE Modifications
The main insight from Table V is that reducing parameters like the bit scale, polynomial modulus degree, or extrema count enables encrypting more values but often comes at the expense of security. For example, decreasing the bit scale from $40$ to $20$ increases the number of encryptable values to $690,000$, but it also lowers the security level. Similarly, a smaller polynomial modulus degree allows for encrypting more values, but it reduces both the security level and the depth of operations. In FHE systems, communication costs are as critical as computational efficiency, especially when encrypted data is transmitted, such as in cloud-based machine learning scenarios.
@@
Bit Scale Polymodulus Degree Extrema Count Encrypted Parameter Count Time (s)
40 32768 8 15000 290
30 16384 8 15000 17
20 8192 8 20000 319
20 8192 1 50000 209
30 8192 1 85000 338
40 4096 1 180000 440
40 8192 1 140000 326
30 8192 1 350000 391
20 4096 1 410000 339
20 32768 1 690000 432
Larger ciphertexts, which result from higher encryption parameters or deeper computational models, require greater bandwidth, increasing communication overhead. As illustrated in Fig. 6, encryption and serialization times grow with matrix size, but serialization dominates at larger scales, highlighting the need for efficient data handling. Thus, optimizing encryption parameters like bit scale or polynomial degree is crucial for balancing security, computational efficiency, and communication costs.
@@ VII-C Efficacy of MQMoE Approach
The dataset by Livingstone and Russo [2018] contains audio-visual recordings of 24 professional actors (12 male and 12 female), each expressing eight emotions: calm, happy, sad, angry, fearful, surprised, disgust, and neutral. The data is synchronized with both speech and song modalities, making it ideal for multimodal emotion recognition tasks that require the joint analysis of audio and visual cues. This richness in data supports the evaluation of the MQMoE approach in handling real-world, emotionally nuanced inputs across different modalities. As shown in Table VII, the findings from the RAVDESS dataset indicate that the MQMoE framework replicates trends observed in prior experiments. Specifically, the application of QFL+FHE combined yields enhanced performance metrics. The incorporation of audio-visual modalities improves recognition accuracy and proves the modularity and adaptability inherent in the MQMoE architecture. This demonstrates its effectiveness in multimodal emotion recognition, presenting a solution that can adapt to multiple modalities.